The width of a rectangle is 5 meters less than its length. The area is 84 square meters. Find the dimensions of the rectangle and the perimeter of the rectangle.
Answer by peter_mccarter(5)

The width of a rectangle is 5 meters less than its length. The area is 84 square meters. Find the dimensions of the rectangle and the perimeter of the rectangle.
First it helps to draw a picture so draw a rectangle
Remember that area of a rectangle =Length times width and the perimeter is P= 2 times the length plus 2 times the width
So A=LW and P=2L+2W
84= area
L=length
W= length -5
L(L-5)=84 distribute the L and you have L2 -5L=84 now subtract 84 from both sides and get L2 –5L-84 next factor to get (L+7)(L-12) next set both to =0 (L+7)=0 (L+12)=0 solve for L
L=-7 and L=12 you can’t have a negative length so -7 is out. So L=12 now plug in to the original area equation
84=LW 84= 12w now divide both sides by 12 to isolate W, this gives you W=7
L=12 and W= 7
For perimeter use 12 for the length and 7 for width Remember that P= 2L+2W
P=2(12) +2(7) =38